### v1

# Towards Human-AI Synergy in UI Design: Enhancing Multi-Agent Based UI Generation with Intent Clarification and Alignment

Abstract.
In automated user interface (UI) design generation, a key challenge is the lack of support for iterative processes, as most systems only focus on end-to-end generation of designs as starting points. This results from (1) limited capabilities to fully interpret user design intent from text or images, and (2) a lack of transparency, which prevents designers from refining intermediate results. To address existing limitations, we introduce PrototypeAgent, a human-centered, multi-agent system for automated UI generation. The core of PrototypeAgent is a theme design agent that clarifies implicit design intent through prompt augmentation, coordinating with specialized sub-agents to generate specific components. Designers interact with the system via an intuitive interface, providing natural language descriptions and layout preferences. During generation, PrototypeAgent enables designers to refine generated intermediate guidance or specific components, ensuring alignment with their intent throughout the generation workflow. Evaluations through experiments and user studies show PrototypeAgent’s effectiveness in producing high-fidelity prototypes that accurately reflect design intent as well as its superiority over baseline models in terms of both quality and diversity.
Interactive design, User interface, Large language models, Image generation, Design Intent Clarification, Multi-Agent, Intent-Design Alignment, Automated Prototype Generation
copyright: nonecopyright: acmcopyrightjournalyear: 2022doi: XXXXXXX.XXXXXXXbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology(UIST’ 24)price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and tools
Figure 1. An overview of the interactive user interface (UI) prototype creation process facilitated by PrototypeAgent, our multi-agent generation system. (A) Users input their design intentions through prompts and initial UI layouts. (B) Using our advanced multi-modal, multi-agent system, we produce detailed prototypes. (C) These prototypes offer a top-down perspective, allowing users to interactively refine content, ranging from themes to specific components. (D) The output is a high-fidelity (both visual and functional) prototype.
1. Introduction
In recent years, the field of user interface (UI) design has seen the emergence of various tools and methods aimed at assisting designers. Notable among these are retrieval-based methods (Chen et al., 2020a; Li et al., 2021; Bunian et al., 2021; Huang et al., 2019) and generative approaches (Cheng et al., 2023; Biswas et al., 2021). These advancements aim to streamline the design process and enhance productivity and creativity.
In designers’ daily creation workflows, they typically rely on professional GUI prototyping tools, such as Sketch (Sketch, 2024), Adobe XD (Adobe, 2024), Figma (Figma, 2024). These typically offer a combination of fundamental GUI components, templates and abundant manual operations. However, they can not automatically generate customized results based on design requirements, which limits their support for the creative process and efficiency of the overall design process.
To support the design process, many retrieval based methods have been proposed to offer inspirations by retrieving relevant UIs given designers’ initial ideas, like sketch, wireframe and natural language descriptions (Chen et al., 2020a; Huang et al., 2019; Li et al., 2021; Bunian et al., 2021; Bernal-Cárdenas et al., 2019; Kolthoff et al., 2023a). Chen et al. (Chen et al., 2020a) leverages wireframes as a bridge to retrieve corresponding high-fidelity functional and visual prototypes. Swire (Huang et al., 2019) learns the distance between designers’ sketch and high-fidelity UI design to enable sketch based search. While Guigle (Bernal-Cárdenas et al., 2019) use rich metadata from UI screenshots and app stores to enable natural language based query. However, these methods often suffer from limitations in terms of visual fidelity, creativity, functional fidelity and reusability. Recent advance in generative methods such as layout2image (Biswas et al., 2021), VAE (Razavi et al., 2019), MidJourney (Midjourney, 2024) and stable diffusion (Rombach et al., 2022b) has increased their use to support creativity, which showcase remarkable creative potential, however, their practical applications often yield unstructured and non-editable outputs. This limits their usefulness in UI design scenarios, where further manipulation and customization are frequently needed to meet designers’ specific requirements. Additionally, small elements like icons, which carry nuanced and meaningful details, can not be handled well by these models, further exacerbating these challenges. Moreover, both retrieved-based and generative methods often require designers to manually reconstruct these UI using professional tools to cater for their needs.
To address these challenges, we observe that industry tools such as Uizard (Uizard, 2024), Vercel’s V0 (Vercel, 2024), and Figma plugins (uxp, 2024) have made progress by using AI to generate initial design concepts, aiming for automated design generation. However, how AI can continuously support designers throughout the entire design iteration workflow, from concept to refinement, remains largely unexplored. Specifically, questions remain about which aspects of the design process should remain under human control, where AI-generated processes require transparency, and how automation can be fine-tuned by humans to maximize its effectiveness throughout the workflow.
To gain deeper insights into the current human-AI collaboration challenges in the UI design domain, we interviewed 10 professional UI/UX designers. This revealed five key shortcomings in the current interaction between humans and AI-driven design tools, including the lack of: (F1) seamless and flexible integration of knowledge bases, (F2) Design modality requirements for accurately expressing functional design intent with AI, (F3) Implicit design intent clarification to enhance design descriptions, (F4) Decoupling the generation process for human-centered intent-design alignment, and finally (F5) Ensuring design coherence throughout the process.
In response to these observations, we introduce PrototypeAgent, a novel interactive design system that enables multi-agent collaboration for UI prototype generation with intent clarification and alignment. This system empowers designers to craft high-fidelity prototypes that are editable, customisable and comprehensible at each stage of the generation process. As depicted in Figure 1, PrototypeAgent begins with two design inputs: a textual description and a wireframe layout (Figure 1 (A)). The text outlines the general design requirements (e.g., “Starting page for PrototypeAgent: an intelligent design assistant”), while the wireframe offers a preliminary UI layout, enabling designers to convey their initial and nuanced design concepts.
In the prototype generation phase (Figure 1(B)), PrototypeAgent employs a top-down strategy, beginning with a central agent that establishes the overarching theme and performs implicit intent clarification through prompt augmentation. Following the decoupling approach, the central agent coordinates the automatic, iterative creation of components, leveraging a cache pool to seamlessly integrate details from various sub-agents, each specially designed with expertise in different aspects. These specialized agents, including a text content generation agent, image content generation agent and retrieval-based icon agent, ensure alignment with the overall design. This integrated approach allows PrototypeAgent to maintain aesthetic consistency, elevating the quality and coherence of the final prototype.
In the review and editing phase (Figure 1 (C)), PrototypeAgent offers deeper insights and detailed control within its interactive environment. It facilitates precise, iterative refinement, balancing automated generation with human customisation for efficient and personalized integration. Additionally, the generated prototypes (Figure 1 (D)) can be saved in SVG or JSON formats, highlighting the system’s practicality and supporting high-quality creative output.
For evaluation, we first conducted two automated quantitative evaluations, which validate PrototypeAgent’s ability to generate realistic and detailed UI designs. Our ablation study further highlights the significant impact and necessity of each module. In addition to the automated evaluations, we carried out two user studies involving 16 participants to assess the perceived usefulness of our work. The findings suggest that PrototypeAgent was positively received and showed significant potential in addressing the five identified challenges. Our work includes three main contributions:
  • To the best of our knowledge, this work is among the first to report findings from semi-structured interviews with professional UI/UX designers working with AI-powered design tools, detailing their current workflows and identifying five key shortcomings.
  • We introduce PrototypeAgent, a novel multi-agent prototype generation system that addresses implicit clarification through design prompt augmentation and ensures intent-design alignment and design coherence by decoupling the generation process. It transforms user descriptions and low-fidelity wireframes into detailed prototypes based on our specialized domain-specific knowledge bases, with user feedback integrated at each stage to ensure precision.
  • We conducted both quantitative and qualitative evaluations to confirmed the effectiveness of our human-in-the-loop prototype generation approach and demonstrated its potential to enhance the UI/UX design process.
2. Related Work
Table 1. Comparison of Design and Generative Tools Used in Professional UI/UX Workflows
Design Tool Generative Tool
Adobe Illustrator Figma / Adobe XD Midjourney / Stable Diffusion Vercel’s V0 Uizard PrototypeAgent (Our)
Task Design Tool Design Tool Image Generation UI Generation UI Generation UI Generation
Target Users Designers Designers Designers / End Users Developers Designers Designers
Input SVG SVG NL description, UI screenshot NL description, UI screenshot NL description, UI screenshot NL description, Wireframe
Output SVG SVG Image (screenshot) Code + Rendered page SVG SVG
Knowledge Base N/A N/A General image datasets Confidential Knowledge Base Confidential Knowledge Base Real-world UI, UI Components, Icon datasets, LLM-generated UI semantic datasets
Implicit Intent Clarification N/A N/A Not Supported Not Supported Not Supported Supported (through the Theme Design Agent prompt augmentation)
Human-Centered Intent-Design Alignment N/A N/A Not Supported Not Supported Not Supported Supported (through generation decoupling and explicit intermediate guidance)
Design Coherence Among Elements (Semantic + Style) Manual Manual Implicit Unknown Unknown Explicit (through decoupling and multi-agent explicit generation)
2.1. GUI Design Tools and Techniques
In professional GUI design, tools like Sketch (Sketch, 2024), Adobe XD (Adobe, 2024), and Figma (Figma, 2024) are popular due to their extensive libraries, templates, and high-fidelity prototyping features. Adobe Illustrator (Adobe, 2024) is preferred for detailed illustrations. However, these tools lack robust automated generation capabilities, which limits efficiency and accessibility for designers with varying expertise. To address these constraints, research in Human-Computer Interaction (HCI) and Software Engineering (SE), along with commercial products, has focused on enhancing automation in design processes.
For creative design inspiration, text-based GUI retrieval, such as Guigle(Bernal-Cárdenas et al., 2019), leverages automated crawling and natural language processing to perform efficient searches through app hierarchies. Systems like Gallery D.C.(Chen et al., 2019), GUI2WiRe(Kolthoff et al., 2020), and RaWi(Kolthoff et al., 2023a) have further enhanced component extraction from screenshots, enabling flexible search based on dimensions, color, and text. Visual and multi-modal retrieval methods extend this by incorporating richer inputs such as screenshots and wireframes. Methods like WAE(Chen et al., 2020a), VINS(Bunian et al., 2021), and Swire(Huang et al., 2019) bridge the gap between low and high-fidelity designs, while Screen2Vec(Li et al., 2021) introduces multimodal embeddings for diverse GUI content. In high-fidelity prototyping, tools like GUIGAN(Zhao et al., 2021) and research by Forrest et al.(Huang et al., 2021) focus on component retrieval and arrangement for detailed prototypes. However, these retrieval-based methods still rely heavily on existing designs, limiting creative flexibility. They address early-stage inspiration but leave much of the manual effort required for design refinement and customization.
On the generative front, techniques like Layout2Image(Biswas et al., 2021), VAE(Bengio and LeCun, 2014), MidJourney(Midjourney, 2024), and Stable Diffusion(Rombach et al., 2022b) offer new possibilities for design generation. However, these methods often result in unstructured, non-editable outputs, making them challenging for GUI design. Projects such as PLay(Cheng et al., 2023) and DocSynth(Biswas et al., 2021) have made contributions to layout generation, facilitating the creation of low-fidelity prototypes. However, they provide limited support for detailed component creation or high-fidelity refinement. Industry tools like Uizard(Uizard, 2024), Vercel’s V0(Vercel, 2024), and Figma plugins(uxp, 2024) provide automated prototype generation, yet they primarily offer starting points. The missing puzzle piece is how AI can continually collaborate with designers throughout the design iteration workflow, from concept to refinement.
Our work addresses these limitations by making Implicit Intent Clarification throughout the generation process. Designer prompts are often vague, and the AI generation process is opaque, making it hard for designers to refine outputs effectively. Clarification is needed in two directions: (1) helping designers provide more precise inputs and (2) making the AI’s decision-making process transparent to allow better control and modification. As reviewed in Section 3 and Table 1, our approach bridges this gap by ensuring intent clarification before the entire generation and decoupling content generation into a more controlled manner.
2.2. Interactive Applications with Large Language Models
Large Language Models (LLMs) have significantly influenced HCI, showcasing remarkable capabilities in zero/few-shot learning and conversational interactions. These models have been instrumental in various applications, enhancing user experience and efficiency. Wang et al.(Wang et al., 2023a) employ LLMs to facilitate versatile conversational interactions within mobile UIs, a notable step in making user interfaces more accessible and intuitive. Stylette(Kim et al., 2022) innovatively integrates LLMs to interpret natural language commands for web design modifications, effectively translating them into corresponding CSS properties.
Further extending the LLMs’ utility, MenuCraft (Kargaran et al., 2023) demonstrates their effectiveness in generating menu designs, while SUGILITE (Li et al., 2017) leverages them for automating tasks on smartphones. Duan et al.(Duan et al., 2024) have explored LLMs’ potential in providing automatic feedback for UI designs based on guidelines, indicating the model’s analytical strength. Similarly, CollabCoder(Gao et al., 2024b) utilizes LLMs for collaborative qualitative analysis, emphasizing their role in team-based settings. Moreover, VIVID (Choi et al., 2024) presents an application of LLMs in generating vicarious dialogues from lecture videos, showcasing their adaptability in educational contexts.
Some previous works have explored agents that clarify intents during the process of human-AI collaboration through interactions between end users and AI using natural language instructions (Li et al., 2018; Srivastava et al., 2017). SOVITE (Li et al., 2020) expanded on this by enhancing system transparency through a mutual disambiguation pattern (Oviatt, 1999), where inputs from one modality help clarify inputs from another, allowing for breakdown repair. However, the interactive design generation process remains under-explored, as it is often unreliable and confusing. Motivated by SOVITE and related insights, and as shown in Table 1, our decoupling generation method with Human-Centered Intent-Design Alignment addresses this gap by allowing both the designer and AI to collaboratively understand and edit intermediate guidance and results, improving clarity and disambiguation throughout the process.
2.3. Large Language Models for Agent
The emergent capabilities of LLMs, such as zero-shot prompting and complex reasoning (Wu et al., 2022a; Vaswani, 2017; Brown, 2020), have spurred their integration into advanced agent research (Betz et al., 2021; Liu et al., 2021; Lu et al., 2021; Wu et al., 2022b). Notable examples include ReAct’s (Yao et al., 2022) use of chain-of-thought prompting (Feng and Chen, 2024) for generating reasoning traces and task-specific actions, and tools like AutoGPT (Gravitas, 2023), Chameleon (Lu et al., 2024), and VOYAGER (Wang et al., 2023b) make use of their internal knowledge bases to automate tasks by formulating a curriculum of multiple subgoals to achieve a high-level goal.
These LLMs have shown exceptional conversational engagement, decision-making, and task-completion abilities, significantly improving communication between users and agents. In the HCI research, applications like AudiLens (Park and Choi, 2023) have developed intelligent agents for public speech practice simulations, and GestureGPT (Zeng et al., 2023) has pioneered a dual-agent gesture understanding framework. Dang et al.(Dang et al., 2023) integrate user instructions into UI variants, enhancing nuanced user-LLM interaction. Gao et al. (Gao et al., 2024a) further identified key phases and interaction modes in Human-LLM interaction.
In the visual tools domain, the integration of LLMs with models like HuggingGPT(Shen et al., 2023) and visual foundation models(Ramesh et al., 2021; AI, 2022, 2023) has marked significant progress. However, there remains a gap in fully leveraging the combined potential of visual models and LLMs for improving design generation, particularly in interactive UI design. Our work addresses this by introducing a divide-and-conquer approach to natural language descriptions as a design-specific Domain-Specific Language (DSL) that is easily interpretable by both humans and LLMs. This DSL allows for detailed descriptions of global and local design elements, enhancing prompt clarity.
This form of prompt augmentation functions as a “software tool,” enabling designers to interact more efficiently with the system. By focusing less on crafting complex prompts and more on the core content of the design prototype, designers can actively participate in the design process. This provides real-time control over intermediate steps, allowing them to adjust the design direction, thereby improving the quality and coherence of the final output.
3. Formative Study and Findings
In this section, we conducted a study to gain insights from a professional designer’s perspective on the current design process when working with AI-powered design tools, identify the challenges UI designers face and explore potential improvements in design tools to better support their workflows.
Utilizing convenience sampling, we conducted interviews with professionals in the industry who have substantial experience in UI/UX design. In total, we interviewed 10 UI/UX designers (2 males, 8 females) from five different companies. These companies range from small startups to large corporations. Our participants had a diverse range of working experience: four had 1-3 years of experience, three had 3-5 years, and three had over five years in the field. Their ages varied from 23 to 34 years, averaging at 28.
The interviews were conducted through online video meetings and lasted approximately 30 minutes on average. These interviews were recorded and transcribed verbatim. The designers were asked ten questions to assess the current status of design tools and identify areas where improvements could be made  [^1] The questions are included in the supplementary materials. Participants were questioned about their design workflow, software tools used and their benefits, the possible application of AI support at different stages, and the investigation of AI tools to streamline repetitive and less technical design tasks. Furthermore, we asked for their assessment of an example UI design generated by a generative model, the strengths and drawbacks of their current design tools, the demand for AI-assisted creation, and their views on aesthetics and colour requirements were also explored. To analyse the data, we adopted a thematic analysis approach (Clarke et al., 2015). For efficiency, the coding process was carried out on a one-interview-one-coder basis. Afterwards, the two coders collaboratively performed aggregation steps. From our thematic analysis, we extracted 96 in-vivo codings. The study results were twofold: first, we summarized the key advantages and disadvantages of the current design and generative tools, as reported by the designers. Second, we identified five primary actionable requirements, elaborated in Section 3.1.
3.1. Findings
Our findings indicated a prevalent use of design search tools/platforms like Google, Dribbble (Dribbble, 2024), and iconfont (Alibaba, 2024), and creative design platforms such as Figma (Figma, 2024), Adobe XD and Adobe Illustrator (Adobe, 2024). Additionally, new AI generation tools such as Midjourney (Midjourney, 2024) and Stable Diffusion (AI, 2022, 2023) have been incorporated into their daily routines. Furthermore, participants also mentioned newly released AI generative design tools, including Uizard (Uizard, 2024) and Vercel’s V0 (Vercel, 2024). However, due to their current beta status, instability, and the requirement of additional subscription fees, these tools have not yet been integrated into the designers’ daily practices. Participants acknowledged the value of these tools in fostering creative ideas and producing well-designed UIs, while also identifying specific areas for enhancement.
Our formative interviews confirmed the limitations of current input and output modalities used by designers, highlighting the need to address design aspects that better support the generation process. Unlike commercial tools, our PrototypeAgent focuses on implicit design intent clarification, intent-design alignment, and ensuring coherence between design elements throughout the prototype generation process. Table 1 in Section 2.1 compares popular design and generative tools used in participant workflows, helping to illustrate our findings and the advantages of our proposed method. We have identified five key findings as follows:
F1. Seamless and Flexible Integration of Knowledge Bases While there are numerous design search tools and generative models available, a design space related to the use of knowledge bases still needs to be considered. Participants in our study highlighted that knowledge bases should be embedded directly into the design workflow, rather than existing as isolated resources. This integration can enhance the accuracy and automation of content generation and use, while also reducing the risks of generating irrelevant or disconnected designs. As one participant remarked, “I regularly discover designs or elements that match what I’m looking for. However, seamlessly integrating them into my design process remains difficult. AI-powered tools should make this easier.”
In addition to seamless integration, participants highlighted the need for systems that support the ongoing expansion of their knowledge bases. They expressed a desire for scalability, allowing designers to build upon and customize their existing resources as their needs evolve. To address this, our approach incorporates a dynamic, non-training method for utilizing the knowledge base: retrieval-augmented generation (RAG). This ensures continuous augmentation of the designer’s knowledge base without requiring extensive retraining, making the system both flexible and adaptive.
F2. Design Interaction Modality Requirements Our findings indicate that design interaction modalities should prioritize controllability in terms of functionality and layout, as well as the editability of outputs, as outlined in Table 1 Input and Output.
Traditional design tools often require manual creation of designs, typically in SVG format. Emerging generative tools offer designers the ability to input natural language descriptions and high or mid-fidelity design images (e.g., screenshots). However, interviews with participants revealed that the ability to precisely and controllably describe their design intent is a crucial requirement for effective input. As one designer expressed, “Text-only or image-only input doesn’t fully capture what I’m imagining, especially regarding the functionality of each component, which limits their practical use.” Participants noted the advantages of combining wireframe screenshots with text descriptions, which not only provide control over functionality and layout but also preserve creativity and the automation benefits of AI-powered generation tools.
Regarding Output, participants emphasized that editability is essential. While SVG format allows for easy modifications, other formats, such as code (e.g., Vercel’s V0), are challenging for designers lacking programming skills. As one participant shared, “Getting a design from MidJourney is amazing, but the real work begins with the day of editing that follows.” This underscores the need for output formats that support easy editing and refinement.
F3. Bridging AI and Designer Intent through Implicit Intent Clarification Designers expressed that AI and designers should collaborate to enhance each other’s work, especially when it comes to clarifying design intent. One of the key areas highlighted is the need for AI tools to assist in refining the initial input provided by designers. Due to unfamiliarity with prompt structures or unclear design objectives, designers often start with simple inputs like “design a landing page” without specifying the granular details of their requirements. AI should work with designers to clarify and refine these inputs.
As one participant noted, “I believe AI understands the kind of instructions it needs better than I do. If it could provide feedback, I could express my design goals more quickly and more accurately.” This emphasizes the importance of automated intent clarification as part of the AI-human collaboration process, aligning with the concept of achieving common ground in communication theory (Clark, 1991).
To address this, our system incorporates an Implicit Intent Clarification step, which refines prompts by regenerate them into specific design aspects. As highlighted in Table 1 under Implicit Intent Clarification, this is an important aspect of AI assistance that is currently missing in other AI generation tools.
F4. Decoupling the Generation Process for Human-Centered Intent-Design Alignment Studies (Dang et al., 2023) have highlighted the lack of transparency and explainability in end-to-end generation methods, particularly in achieving intent-design alignment from themes to local components. Designers often need to make adjustments—changing design styles, editing text, or replacing images—but the opacity of these methods forces them to “guess” the correct input phase, wasting time and reducing accuracy.
As shown in Table 1 under Human-Centered Intent-Design Alignment, current tools lack clear explanations of their generation steps. Designers called for more human-centered mechanisms to better understand and refine the process. One participant remarked, “If AI tools could reveal their thought processes and allow editing, it would greatly simplify tailoring the results to our needs.”
Our approach addresses this by decoupling the generation process and integrating intent-design alignment. This allows designers and AI to collaboratively refine intermediate guidance and results. Using natural language as DSL for global and local descriptions, our method promotes clarity, reduces ambiguity, and supports both large-scale and component-level alignment.
F5. Ensuring Design Coherence Through Generation Processes Maintaining consistency in design elements—such as color, style, and aesthetics—across a project is often a repetitive and labor-intensive task. As shown in Table 1 under Design Coherence Among Elements (Semantic + Style), traditional design tools require manual intervention to achieve visual cohesion. While newer generative tools can use text descriptions or reference images to assist with this, their ability to maintain consistency remains unclear. As one participant noted, “Creating a unified look for a new app involves a lot of repetitive work.” This underscores the need for tools that can automatically maintain design consistency.
Our approach addresses this challenge by employing an agent as a central controller that manages design consistency across various components. It not only ensures coherence but also explicitly presents the natural language content describing design consistency, allowing designers to interact with and refine the results efficiently.
4. Approach
4.1. Overview
To improve the findings identified in Section 3 and actualize the design space outlined, we introduce PrototypeAgent: a system that harnesses multi-agent collaboration for the explainable generation of UI prototypes. This system aims to enhance mutual disambiguation in Human-AI Collaboration for creating high-fidelity prototypes by offering a decoupled generation process.
While large language models excel in contextual understanding and response generation, UI generation poses unique challenges that require deep domain knowledge to ensure high-quality design. A simple end-to-end approach may fall short of capturing the complexities of UI design intent. Therefore, we adopt a divide-and-conquer, top-down strategy that decouples the generation process into sub-steps. A central managing agent oversees these steps to ensure coherence across the design. By using natural language descriptions as a DSL, our method enables designers and AI to collaboratively understand and refine intermediate results.
As depicted in Fig. 2, PrototypeAgent integrates knowledge bases with four specialized agents: Theme Design Agent $A_{theme}$, Text Content Agent $A_{text}$, Image Content Agent $A_{img}$, and Icon Agent $A_{icon}$. We first collected two knowledge bases offline to import relevant design knowledge, corresponding to F1. Seamless and Flexible Integration of Knowledge Bases in the Section 3.1. The online system commences its operation by processing design inputs, including a prompt and UI layout, as specified in F2. Design Interaction Modality Requirements. The Theme Design Agent initiates the process by generating a high-level theme description for F3. Implicit Intent Clarification and corresponding image, setting the stage for a consistent design narrative in line with F5. Ensuring Design Coherence.
Functioning as the primary director, $A_{theme}$ delegates specialized tasks to $A_{text}$, $A_{img}$, and $A_{icon}$ for the detailed creation of each design component, addressing F4. Human-Centered Intent-Design Alignment and F5 respectively. The effective coordination between the central and supporting agents, facilitated by a cache pool, is crucial for maintaining the design’s contextual integrity.
This methodology fosters clarity, disambiguation and iterative interactions, allowing PrototypeAgent to progressively refine the design. Designers are thus provided with an environment that not only generates designs but also supports iterative refinements to align closely with their intent.
4.2. Knowledge Base Construction
Employing domain-specific knowledge can harness the creativity of Large Language Models (LLMs) while enhancing the quality of generation (Wang et al., 2023b; Lu et al., 2024). We collected two kinds of knowledge, namely UI knowledge (pairs of layout/theme descriptions with local component descriptions) and Icon Knowledge (Pairs of icon SVG code with semantic descriptions) for $A_{theme}$ and $A_{icon}$.
Table 2. Blip2 model’s VQA instruction templates for RICO screenshots
Theme Design Attributes Blip2 Instruction Templates
Theme Color ¡Image¿ “Question: What is the background color of this screenshot? Answer:”
Primary Color ¡Image¿ “Question: Besides the background, what’s the dominant color in this image? Answer:”
Theme Description ¡Image¿ “Question: Can you describe this screenshot in detail? Answer:”
App Category ¡Image¿ “Question: Which category does this app belong to? Answer:”
4.2.1. UI Knowledge Base
We build our knowledge base regarding the UI composition and semantic knowledge by considering two datasets (Rico (Deka et al., 2017) and Screen2words (Wang et al., 2021)) and using one large language model, Blip2 (Li et al., 2023). The aim of this module is to obtain two kinds of knowledge: (1) UI Composition Knowledge: ¡component types¿¡bounding boxes¿; (2) UI Semantic Knowledge: ¡text content/icon descriptions¿¡high level description¿¡theme design description¿. An example can be seen in Fig. 4(c).
1) UI Composition Knowledge. Rico Dataset (Deka et al., 2017) is one of the most comprehensive open-source UI datasets available, which contains around 66k distinct UIs from over 9.7k Android apps across 27 categories. This dataset includes the UI screenshots and their view hierarchy information, which expose UI elements used, their attributes like text, bounds and class, and the composition of these UI elements. We extracted class and bounds from these metadata, and formed the UI composition knowledge (¡component types¿¡bounding boxes¿) for each UI.
2) UI Semantic Knowledge. To obtain the UI semantic knowledge, we utilize the Rico dataset, Screen2Words and a visual question-answering model to obtain the fine-grained component descriptions, high-level UI descriptions and theme design descriptions, respectively. The fine-grained component description can be obtained by parsing the text, content-description from the UI metadata from the Rico dataset. Thus, we obtained ¡text content/icon descriptions¿.
While Rico contains the metadata of the composition and text contents of the UI, it lacks the high-level UI description. We obtained this data through Screen2Words (Wang et al., 2021), which augments the Rico dataset by hiring crowdsourcing workers to provide  112k high-level textual descriptions for its 22k UI screenshots. Through this dataset, we obtained ¡UI description¿.
Beyond UI functionality semantics, design generation necessitates thoughtful consideration of themes, colours, and the target audience. We adopt Blip2 (Li et al., 2023), which is a zero-shot visual language model, to generate theme descriptions via a visual question-answering approach. We identify four key attributes for theme design: theme colour, primary colour, theme description, and app category in Section 3. For each attribute, we create a specific question, pairing it with a UI image, and then input it into Blip2 to derive the answer. The questions used to extract these attributes can be found in Table 2 Finally, these three descriptions are concatenated together, and form UI Semantic Knowledge: ¡text content/icon descriptions¿¡ high-level description¿¡theme design description¿.
Figure 2. Overview of PrototypeAgent System: This figure illustrates the main process of PrototypeAgent in response to a designer’s input and UI wireframe. PrototypeAgent utilizes a multi-agent approach for the interactive generation of UI prototypes. It encompasses four specialized agents—Theme Design $A_{theme}$, Textual Content $A_{text}$, Image Content $A_{img}$, and Icon $A_{icon}$. The Theme Design Agent $A_{theme}$ acts as the central coordinator, steering the collaborative efforts of the three sub-agents. By leveraging a cache pool, PrototypeAgent adeptly integrates the contributions from each agent to ensure a cohesive alignment with the overall design context. This process not only generates accurate prototypes but also provides explainable intermediate results, enabling designers to conduct thorough reviews and make precise edits.
Figure 3. Examples of SVG code renderings for icons and their corresponding semantic descriptions
4.2.2. Icon Knowledge Base
As visual shortcuts, icons improve the user experience in UI design. They enable intuitive navigation and enhance the visual appeal. To enhance a great generation capability and enable designers to modify based on their needs, we collected an icon knowledge base in SVG format for Icon Agent. This choice not only ensures compatibility across modern browsers but also guarantees that icons maintain their clarity when scaled. The designer can also modify the colour, style, and shape of the icons based on their specific requirements.
We collected the data from Google Material Design Icons (Google, 2024), a high-quality repository that stores over 900 diverse icons (in SVG format and with text description). These symbols adeptly convey universally recognized actions or objects, which would be ideal for design generation. Fig. 3 shows some of the collected icon examples with their corresponding semantic descriptions, such as “add shopping cart”, “alarm” and “bookmark”, to illustrate the semantic usage of these graphs. Therefore, we obtained our Icon Knowledge Base: ¡icon SVG code¿¡semantic description¿.
4.3. Theme Design Agent
The Theme Design Agent functions as the central supervisor in the UI design process, leveraging domain knowledge from the UI knowledge base (Section 4.2). Its primary role is to perform implicit intent clarification through design prompt augmentation, setting the overall style and orchestrating the generation of a global theme description. By ensuring design coherence through the corresponding theme image, this dual-modality approach guarantees a high-quality UI design with consistency across all elements. The agent also collaborates with designers to align and refine design descriptions, facilitating a seamless design process.
Moreover, the Theme Design Agent works in tandem with a cache pool, collaborating with other agents to refine and detail specific components of the UI. The operational phases of the Theme Design Agent comprise four key stages: (1) Knowledge Retrieval: Accessing and utilizing domain-specific information. (2) Theme Description Generation: Facilitating implicit intent clarification and crafting a comprehensive and cohesive theme description. (3) Theme Image Generation: Producing a visual representation of the theme. (4) Sub-agent Execution: Coordinating sub-agents to execute tasks, ensuring intent-design alignment throughout the generation of each component. These stages are elaborated upon in subsequent sections, offering a detailed insight into each phase of the theme design process.
4.3.1. Knowledge Retrieval
Domain-specific knowledge enhances the accuracy of LLM-generated content. Recognizing the LLM’s token input limitations and the complexities of fine-tuning, we introduce a knowledge retrieval phase, infusing external, domain-specific knowledge into our generation process.
Based on the design prompt ($In_{p}$) and UI layout ($In_{l}$), which is provided in a bounding box format accompanied by component labels as an example depicted in Fig. 4(B), we want to retrieve the most relevant knowledge from our large knowledge base. To do so, we concatenate these two information together, and encode them into one latent vector $\textbf{Emb}(In)$ as the query vector, where $In=In_{p}+In_{l}$. We use the TEXT-EMBEDDING-ADA-002 embedding model (OpenAi, 2024a)). Similarly, we also embed each piece of UI knowledge related to UI Composition and Semantic Understanding into a latent vector ($\textbf{Emb}(kb_{j})$) using the same embedding model. After that, we compute the cosine distance between the query vector and each knowledge vector, and retrieve the top-k results to instruct our multi-agent system. We denote the retrieved knowledge as $refer_{i}$.
Our preliminary experiments, alongside findings by Wang et al. (Wang et al., 2023a), indicate that when employing related knowledge as few-shot prompting, the initial example tends to be the most influential. Subsequent examples often provide diminishing returns in focusing the model’s output. Furthermore, given the input length limitations of language models, which restrict the number of exemplars in the prompt, we set the number of references to 2 (i.e., $k=2$).
Figure 4. The prompt design for generating theme description. It consists of four parts: (A) Design Prompt, (B) Wireframe Layout, (C) Top-k Retrieved Knowledge Items, and (D) Theme Description Template. In the wireframe layout preview, different colours denote various component types as defined by Enrico (Leiva et al., 2020).
4.3.2. Theme Description Generation
The Theme Description Generation is a key part of our system, leveraging natural language descriptions as a DSL to facilitate implicit intent clarification and bridge the gap between the designer’s intent and the generated results, ensuring coherent design narratives.
To illustrate the mechanisms driving this process, we detail a structured expression: Given a design prompt $In_{p}$, a UI wireframe layout $In_{l}$, and the top $k$ retrieved knowledge items $\{\sum_{i=0}^{k}refer_{i}\}$ (where, $k=2$), these components are concatenated with the system’s theme description prompt $P_{theme}$ to formulate a comprehensive input for our agent. Formally, the amalgamated prompt $P$ is delineated as:
$P=In_{p}\oplus In_{l}\oplus\left(\sum_{i=0}^{k-1}refer_{i}\right)\oplus\left(P_{theme}\right)$
where $\oplus$ symbolizes concatenation, and the summation symbol $\sum$ here specifically indicates the sequential concatenation of retrieved knowledge items. With this integrated input at its disposal, the Theme Design Agent commences the generation of the Theme Description.
Upon completing the theme description generation, the resultant theme description is denoted by $Res_{theme}$. We show an example in Fig. 4, where the design prompt $In_{p}$ is depicted in part (A), the wireframe layout $In_{l}$ in part (B), the top $k$ retrieved knowledge items $\{\sum_{i=0}^{k}refer_{i}\}$ in part (C), and the system’s theme description prompt $P_{theme}$ is presented in part (D).
4.3.3. Theme Image Generation
The objective of generating a theme image is to visually guide the overall design and ensure intent alignment. allowing the sub-agent to generate coherent and consistent design.
We consider the Stable Diffusion model (Rombach et al., 2022a), a state-of-the-art text-to-image generation, as our main model. The main assumption of the stable diffusion model is that given a random image, we can gradually denoise it to ultimately obtain the meaningful UI design that we want. Therefore, the primary objective of stable diffusion is to predict and progressively eliminate noise from the initial image. However, as this model only considers the text condition, it suffers from limited control over the spatial composition of the image, a crucial aspect of our UI design generation. To address this, we integrate ControlNet (Zhang et al., 2023), which augments the diffusion model by providing enhanced spatial control over each module. In detail, the stable diffusion model contains three parts: Text Encoder, Denoising Module (i.e., UNet) and Autoencoder Decoder.
Figure 5. Layout-guided text-to-image Generation.
The Text Encoder encodes the design prompt, Denoising Module denoises the original random images in several rounds, and the Autoencoder Decoder finally generates the image. ControlNet controls the generation by manipulating the denoising module by inserting an additional spatial condition to UNet. We employ UI layout as the spatial condition.
Specifically, as seen in Fig. 5, this theme image generation module takes three inputs: (a) design prompt (the text condition);(b) a latent image generated from Gaussian noise $\sim N(0,1)$ (the original image), and (c) UI layout (the spatial condition). We first put (a) design prompt into the text encoder of diffusion model, pre-trained CLIP ViT-L/14 (Frans et al., 2022) text encoder, and feed (c) UI layout to ControlNet, to obtain their embeddings respectively. We then gradually denoise (b) the latent image through the Denoising Module with the control from ControlNet. Finally, we obtain the output by feeding the denoised image into the Autoencoder Decoder.
In addition, as the stable diffusion model is originally trained on LAION-5B, (Schuhmann et al., 2022) and faces obstacles when generating UI images, which requires a different domain knowledge from general images (Chen et al., 2020b), We finetuned the model using the datasets of UI screenshots and their complementary high-level descriptions collected in Section 4.2.
4.3.4. Sub-agent Execution
During the sub-agent execution phase, the Theme Design Agent identifies the optimal sub-agent corresponding to the component type. We considered 13 component types, as detailed by the Rico dataset. To elaborate, the Text Agent handles “Text Button” and “Text” components. The Image Agent is entrusted with “Image” and “Background Image” components, and the Icon Agent focuses on “Icon” components. For other component types, we seek to render them editable, drawing insights from RaWi (Kolthoff et al., 2023b). The colour for any component is determined by identifying the dominant RGB colour from the image region’s histogram and then representing it in HTML code.
The dynamic between the central agent and the sub-agents is essential to our system’s functionality. When a central agent engages a sub-agent, it uses a combination of past results and a cache pool to inform the sub-agents prompts:
(1) $Cache_{t}=Res_{t-1}+Cache_{t-1}$
(2) $p_{t+1}=p_{sub}+Cache_{t}$
In this context, $Res_{t-1}$ is the output from the sub-agent for the $t-1^{th}$ component. The $Cache_{t}$ represents a cache pool that integrates the previous result with accumulated knowledge from earlier iterations. This cache pool serves as an essential memory function, retaining the design context and facilitating multi-round interactions by allowing the system to “remember” past interactions and decisions. Meanwhile, $p_{t+1}$ functions as the prompt for the $t-1^{th}$ component’s sub-agent interaction, incorporating both the specific prompt $p_{sub}$ for the current sub-agent and the cumulative knowledge in $Cache_{t}$. This mechanism ensures that each sub-agent’s action is informed by the historical context, enabling consistent design, intent alignment and coherent multi-conversation interaction.
Further details on how these prompts are formulated for specific sub-agents, such as the Text Content Agent and Icon Content Agent, are outlined in Sections 4.4 and 4.6, respectively.
4.4. Text Content Agent
The primary role of Text Content Agent is to generate textual information tailored to specific GUI components. We use GPT-4 (OpenAi, 2024b). The system prompt for this agent, represented as $p_{text}$, is: “Based on the theme description and relevant details, provide a text content recommendation for the designated position at [bbox].” In alignment with equation 2, the execution prompt of the Text Content Agent obtains its value from the central agent’s cache, denoted as $Cache_{t-1}$, and is subsequently concatenated with $p_{text}$, to ensure consistency in the system. This systematic approach ensures that the system consistently produces text content that seamlessly integrates with the overall theme of the GUI component.
4.5. Image Content Agent
To enhance the generation quality of local image-associated components and maintain the consistency of the generated outcomes, we also deploy our adaptively fine-tuned stable diffusion model in the Image Content Agent $A_{img}$. We reuse the Stable Diffusion model finetuned in Section 4.3.3 but disable the ControlNet module. Rather than using the latent image generated from Gaussian noise, we extract the area of the image component from the theme image as the input (b). In addition, we use the image description from the generated theme description as the design prompt (a). By harmonizing both textual and visual signals, we can guarantee that the produced content aligns seamlessly with the primary theme design intent.
4.6. Icon Agent
The Icon Agent is crucial for selecting appropriate icons and integrating them into the graphical user interface components. In addition to acting as intuitive visual cues, well-designed icons can improve comprehension and the overall user experience. The system prompt for the Icon Agent, $p_{icon}$, is: “In reference to relevant information and taking into account its positioning at [bbox], and based on the theme description, propose an indicative phrase like “msg” for the “Icon”. As shown in equation 2, the Icon Agent’s execution prompt is based on the central agent’s cache, $Cache_{t-1}$, combined with $p_{icon}$. This approach ensures the icons selected match the GUI design semantically and visually. The Icon Agent then retrieves the optimal icon SVG code from the knowledge base, corresponding to the generated semantic phrase.
Figure 6. Interactive Interface of the PrototypeAgent Online Prototyping Editor
5. Implementation
In the implementation of the interface, we developed a web-based rapid prototyping editor using HTML5 and JavaScript, to offer a live preview of components and an interactive editing environment. Motivated by the findings from Section 3, PrototypeAgent is adept at processing design prompts and wireframe inputs to respond to F2. The Online Editor is designed for easy input, with an editable text box that let designers create task prompts seamlessly.
Adding and configuring UI elements is made simple: designers select an element type and can then resize or reposition it using a convenient drag-and-drop interface, as shown in Figure 6. The editor is structured into distinct parts for optimal user experience: Part A allows for the input of design descriptions, Part B serves as the user selection and edit panel where components such as “Web View” and “Button” can be chosen, and layout actions like “Create Layout” or “Load Layout” are available. Manual adjustments to layout size and positioning are done in Section C, while Section D provides a real-time preview of both the layout and the generated results.
Sequentially, in response to F3 and F5, activating the ”Theme Design” function, the generated theme descriptions are displayed in Section E, and theme image are displayed in Section D. This enhances transparency and explainability in the Human-AI collaboration process, enabling mutual disambiguation. Hitting the “Prototype Generation” button makes our system produce a detailed prototype in Section D, with the components’ explainable contents displayed in Section F, aligned with F4. Designers can then save their prototype with the “Save Prototype” button, exporting it as a JSON file for further use.
Beyond Section D’s preview feature, every part of our system is designed to be editable, empowering designers to efficiently refine and produce their final prototype. Upon generating components via our multi-agent system, designers can refine both the intermediate and final outputs. The system allows for real-time visualization of the application design, and any changes made by the designer are instantly updated in the system with the “Update Edit Result” button, which simplifies group text modifications and streamlines the design refinement process.
For the LLM model’s deployment, we utilized OpenAI’s gpt-4 (OpenAi, 2024b) APIs for textual generation within our multi-agent system and paired it with the text-embedding-ada-002 (OpenAi, 2024a) API for text embedding. All temperature parameters were set to 0.
Regarding the image generation model, we used the runwayml stable-diffusion-1-5 (AI, 2022) checkpoint from HuggingFace to finetune the model, and the Rico dataset. The dataset was divided into training, validation, and testing sets, consisting of 157,123 UIs, 2,364 UIs, and 4,310 UIs, respectively, following Rico (Deka et al., 2017). Training images were adjusted to $512\times 512$ resolutions, and optimization was achieved using the AdamW algorithm with a learning rate set at 1e-5 and a batch size of 1. This training utilized an Nvidia 3090 GPU 24GB VRAM.
6. Evaluation
In this section, we evaluate the performance of generated prototypes. We consider two research questions (RQs):
  • RQ1: How does our approach perform against existing models in terms of quality and diversity of the generated UI design?
  • RQ2: How do PrototypeAgent’s individual modules influence its performance in quality and diversity?
6.1. Evaluation Data
From Section 4.2.1, in total, we collected a set of 3,738 UI textual descriptions, their corresponding wireframes and high-fidelity UI design screenshots. For validation purposes, the corresponding UI screenshots are treated as the ground truth, and we experimented on the test set. Images are resized to dimensions of $512\times 512$.
6.2. Metrics
To assess the quality and diversity of generated UI design, we utilize two metrics: Fréchet Inception Distance (FID) (Heusel et al., 2017) and Generation Diversity (GD) (Cao et al., 2019), which are commonly used in the image generation task (Cao et al., 2019; Ge et al., 2020).
Fréchet Inception Distance (FID) (Heusel et al., 2017) serves to quantify how closely the generated images resemble real ones. This metric computes the statistical difference between distributions of generated images and their real counterparts. Specifically, the FID score is defined as:
(3) $FID=||\mu_{r}-\mu_{g}||_{2}^{2}+Tr(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{1/2})$
Here, $\mu_{r}$ and $\mu_{g}$ denote the mean values of the 2048-dimensional activations of the Inception-v3 pool3 layer for real ($r$) and generated samples ($g$), respectively. Meanwhile, $\Sigma_{r}$ and $\Sigma_{g}$ represent their respective covariances. A lower FID score suggests that the generated images’ distribution more closely matches that of the real images, indicating superior quality and diversity. To compute the FID score, an equal number of real and generated images are fed into the Inception-v3 (Inception, 2015) network.
Generation Diversity (GD) (Cao et al., 2019) can evaluate the diversity of generated prototypes. It calculates the pairwise distances between different UI designs within a generated set. Utilizing Perceptual Hashing (Zauner, 2010), the metric computes these distances, with larger average distances indicating a broader variety of designs within the generated set. The formula of GD is:
(4) $GD=\frac{1}{N(N-1)}\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}d(F_{i},F_{j})$
where $N$ is the total number of UI designs in the generated set, $F_{i}$ and $F_{j}$ represent the feature vectors of the $i$-th and $j$-th UI design, respectively, and $d(\cdot)$ denotes the Euclidean distance. A higher GD means the generated images are more diverse. For a consistent quantitative analysis, we use the same Inception-v3 model to extract features for both FID and GD evaluations.
6.3. RQ1: Comparisons to existing image generation models
6.3.1. Baselines
We consider two state-of-the-art image generation models: stable-diffusion-1-5 (AI, 2022) and stable-diffusion-2-1 (AI, 2023). Stable-diffusion-1-5, released in October 2022, is a widely accepted and stable version of the model. The subsequent version, 2-1, was unveiled in December 2022. It enhances the generation of images with greater diversity and realism, particularly for people, designs, and wildlife. Furthermore, it offers support for non-standard resolution generation. As these two baselines do not incorporate ControlNet module, we consider variants of both by integrating ControlNet, denoted as stable-diffusion-1-5 (with ControlNet), stable-diffusion-2-1 (with ControlNet). In addition, we also employ an ablated version of our approach, PrototypeAgent (w/o ControlNet) as a baseline.
6.3.2. Results
As seen in Table 4, several noteworthy observations can be made:
Dominance in Quality. Our model, PrototypeAgent, consistently outperforms both baseline models in terms of FID scores, regardless of whether ControlNet is utilized. Specifically, a lower FID score suggests that the distribution of generated images more closely matches that of real images. This indicates that PrototypeAgent’s outputs are more realistic, evident from its significantly reduced FID scores: 33.08 without ControlNet and an even lower 23.76 with ControlNet.
Consistent Diversity with Details. GD scores reflect the model’s ability to produce varied yet detailed UI designs. A higher GD indicates more detail, suggesting that the designs are diverse and intricate in their presentation. Without ControlNet, PrototypeAgent achieves a GD of 15.95, showcasing a balanced performance between quality and detailed variety. When integrated with ControlNet, the model achieves a commendable GD of 13.98. Although slightly reduced, this score underscores PrototypeAgent’s ability to produce diverse and detailed outputs, even within layout constraints. Notably, PrototypeAgent’s GD scores, both with and without ControlNet, surpass those of the baseline models. This increase in GD values emphasizes our model’s superior capability in producing designs that are varied and enriched with details compared to its peers.
Impact of ControlNet. Incorporating ControlNet results in noticeable improvements in FID scores for all models. For PrototypeAgent, the FID shows an enhancement of 10.68, representing a substantial 32% improvement. However, the slight decrease in GD, from 15.95 to 13.98, suggests that while ControlNet enhances image realism, it might limit the detail in generative diversity. This trade-off between quality and diversity is anticipated since layout constraints naturally reduce the range of potential outputs.
Comparing Our Fine-tuned Model with Baselines. Contrasted with the baseline models, the advantages of our fine-tuned model become clear. In both scenarios, with and without ControlNet, PrototypeAgent achieves superior FID scores, highlighting its excellence in UI design generation. The consistent GD scores, even surpassing some baselines, confirm the model’s capacity to generate designs that are of high quality and rich in detail. We will provide qualitative evaluation through a user study in Section 7.1.
Answer to RQ1: Our PrototypeAgent outperforms the baseline models in terms of both quality and diversity. The addition of ControlNet optimizes this performance further by introducing layout constraints, reinforcing its potential for generating realistic and detail-oriented UI designs.
6.4. RQ2: Ablation Study
6.4.1. Baselines and Ablation Strategy
To better understand the interaction and individual impact of the decoupled generation mechanisms of PrototypeAgent, we perform an ablation study, sequentially removing each module and evaluating the effect. As seen in Table 4, we carefully crafted six ablations. These modules span across the four systematic phases, namely Knowledge Retrieval, Theme Description Generation, Theme Image Generation, and Sub-agent Execution. The Sub-agent Execution phase, being more granular, was further subdivided to examine the effects of the Text Content Agent, Image Content Agent, and Icon Content Agent individually.
6.4.2. Results
By studying the performance impact when a specific module is absent, we can measure its individual contribution. For instance, removing the Knowledge Retrieval module lets us understand the contribution of our knowledge base in shaping the generated UI designs. Furthermore, these baselines serve as a means to pinpoint the robustness of our model. A model that exhibits minimal degradation in performance across different scenarios showcases its robustness and flexibility.
Figure 7. Generated GUI Examples from Input Layout and Prompt: Corresponding Rico Screenshots, stable-diffusion-1-5 (with ControlNet), stable-diffusion-2-1 (with ControlNet), and Our PrototypeAgent.
Knowledge Retrieval. Fundamental to PrototypeAgent, this module imports essential data from our knowledge base. Excluding it results in an increase of 18.8 in the FID score (from 23.76 to 42.56). This represents the most significant decline in generation quality, underscoring that our knowledge base is pivotal in laying the foundation for high-quality design outputs.
Theme Description Generation. At the heart of generating nuanced and relatable user interfaces, this module crafts detailed textual narratives that align seamlessly with the intended theme. The act of removing it results in an FID score of 28.43 and a GD of 11.77. The degradation of FID and GD scores distinctly shows that the module plays an indispensable role in maintaining design excellence. The Theme Description Generation acts as a directive force, ensuring that the generated outputs are not only contextually meaningful but also visually relevant. Without this narrative guide, the designs might lack depth and context, affecting the overall user experience.
Theme Image Generation. This module goes beyond mere aesthetics and strives to bridge textual theme narratives with consistent visual image. It’s not just about generating visuals but ensuring they are in perfect harmony with the overall design theme. The spike in the FID score to 33.08, when this module is excluded, accentuates its paramount significance. When this module is excluded, the FID score soars to 33.08, emphasizing its important meaning. Beyond the numbers, this increase implies that while text description lays the groundwork, it’s the visual representations that bring them to reality, enriching the design with visual context. Thus, Theme Image Generation, just after Knowledge Retrieval, stands as a cornerstone in influencing the authenticity and relevance of the generated designs.
Sub-agent Execution. This phase ensures that suitable sub-agents are selected for collaboration, optimizing each design component’s realization. Based on the evaluation of the Rico dataset, these sub-agents have specialized roles: the Text Agent managed “Text Button” and ”Text” components; the Image Agent was responsible for ”Image” and ”Background Image”; while the Icon Agent was tailored for ”Icon” components. The slight variations in FID scores — 24.06(+0.3), 24.71(+0.95), and 24.32(+0.56) — when these agents were individually excluded, it became evident from their collective contributions that our fine-grained generation approach is key to driving the model’s peak performance.
Answer to RQ2: Every module in PrototypeAgent plays a crucial role in ensuring superior design quality and diversity. The “Knowledge Retrieval” and “Theme Image Generation” modules are the most influential modules. Meanwhile, the “Theme Description Generation” and various “Sub-agents” work in concert to fine-tune and enrich the final output, adding layers of complexity and refinement to the overall design. This experiment confirms the effectiveness of our decoupling strategy, where specialized agents handle distinct components of the process.
Table 3. Comparative analysis of FID and GD scores among various models with and without the ControlNet. A lower/higher FID/GD means the generated images are more realistic/diverse.
Model FID↓ GD↑
stable-diffusion-1-5 (w/o ControlNet) 69.48 15.93
stable-diffusion-2-1 (w/o ControlNet) 67.15 15.42
PrototypeAgent (w/o ControlNet) 33.08 15.95
stable-diffusion-1-5 (with ControlNet) 54.42 11.48
stable-diffusion-2-1 (with ControlNet) 57.23 11.14
PrototypeAgent 23.76 13.98
Table 4. Results of the ablation study for the different modules in PrototypeAgent.
Model FID↓ GD↑
PrototypeAgent 23.76 13.98
-Retrieved Knowledge Items 42.56 12.14
-Theme Description Generation 28.43 11.77
-Theme Image Generation 33.08 12.95
-Text Content Agent 24.06 13.78
-Image Content Agent 24.71 13.38
-Icon Content Agent 24.32 13.41
7. User Study
To further evaluate the perceived usefulness (i.e., performance and usage) of our system, we carried out two user studies. The performance study focuses on the generated UI design compared to image generation models, and another usage study assesses the usage of the system compared to the latest industrial tools. We aim to answer the following research questions:
  • RQ3 (Performance): What is the perceived satisfaction with PrototypeAgent compared to existing image generation tools?
  • RQ4 (Usage): What is the perceived usefulness of our PrototypeAgent compared to the state-of-the-art industrial tools in terms of the five findings identified in Section 3?
7.0.1. Participants
For both user studies, We recruited the participants through alumni network and partnerships. As a result, we obtained 16 UI/UX practitioners (6 males and 10 females) from various corporate entities to attend our user study. The working experience of these professionals varies, allowing for a broad insight into industry practices and expectations. We had 8 participants with 1-3 years of experience, 5 with 3-5 years, and a further 3 who have been active in the field for more than 5 years. These practitioners come from a spectrum of roles within the UI/UX domain, including UI/UX Designers, UX Engineers, Interaction Designers, and UX Researchers.
7.1. User Study 1: Performance Study
To assess designer satisfaction with the quality of our generated results compared to the baseline model, we conducted a performance evaluation focusing on three key aspects: functional semantics, design aesthetics, and color harmony. These criteria were selected to measure the effectiveness and appeal of the generated designs, ensuring both functional accuracy and visual coherence.
7.1.1. Procedure
Building on established GUI and image evaluation methods (Reinecke et al., 2013; Zhang et al., 2008), participants assessed the quality of mobile GUI designs based on three metrics: functional semantics, design aesthetics, and color harmony. Functional semantics evaluates the relevance and clarity of the generated content, focusing on how well it aligns with the intended functions and the quality of its meaning. Design aesthetics assessed visual appeal (Zhao et al., 2021), and color harmony examined the effectiveness of color combinations. Each design was rated on a 5-point Likert scale. For practical evaluation, 20 design tasks from the Screen2Word dataset (Wang et al., 2021) were selected, and three prototypes were generated per task using both PrototypeAgent and stable diffusion models for comparison. Participants were briefed on these metrics before evaluating the generated GUI designs. They independently scored each design, with the source model of each design concealed to maintain objectivity.
7.1.2. Results & Discussion
As shown in Fig. 8, the GUI designs generated by our model outperformed those generated by other methods, achieving significantly higher scores in terms of design-prototype consistency (Mean=4.13), design aesthetics (Mean=3.83) and colour harmony (Mean=3.62).
Analysis of Functional Semantics. Through detailed analysis of the experimental results, we identified common characteristics of low-scoring prototype designs in terms of functional semantics, such as incomplete structures, basic content, overly small or abrupt images, and an excessive number of components. In contrast, high-scoring prototype designs had a clean layout, moderately rich content, and compatible images. The balance between content richness and layout simplicity was also highlighted as an important consideration.
Furthermore, we found that most of our results demonstrated accurate semantic alignment with the design intent, due to our decoupled generation mechanism. This was evident from the high score of 4.32. Comparatively, this score was significantly higher than that of the stable-diffusion-1-5 and stable-diffusion-2-1 generated results 1.58 and 1.7, yet just 0.19 score lower than the real screenshots (4.51). As demonstrated in Figure 7, the stable-diffusion models largely presented images lacking specific semantic content, thus appearing blurry.
Analysis of Design Aesthetics and Color Harmony. Interestingly, a few of the prototypes generated by our model scored higher than real-world prototypes - a notable accomplishment when compared against the benchmark of real screenshots. In Figure 8 (b), our model surpassed the average score for real prototypes in terms of design aesthetics (3.83 compared to 3.71) and in Figure 8 (c) was slightly lower with respect to colour harmony (3.62 compared to 3.74). This is a significant improvement over the stable-diffusion models. Our generated prototypes closely mirrored real-world prototypes’ overall aesthetics and colour harmony and, in some instances, were of superior quality to poorly designed real-world prototypes.
Analysis of Failure Case. Upon cross-verification of our model’s results against real images, we found some minor errors of our results, such as the over-generation of text leading to typography issues. Participant feedback suggested the further alignment of our model with typography and other design guidelines could address this issue, enhancing the generation performance.
Answer to RQ3: The user study underscored the efficacy of PrototypeAgent in creating GUI designs that excel in design-prototype consistency and aesthetic appeal and colour harmony compared to the state-of-the-art image generation models. Notably, the outputs were found to align closely with real-world screenshots, even outperforming them in terms of design aesthetics. Further analysis underscored that our PrototypeAgent decoupled generation mechanism achieved accurate semantic alignment with the design intent.
Figure 8. Comparative Evaluation of Prototype Generation. “Real” refers to results obtained from corresponding real screenshot images, “s-d-1-5” represents the outcomes of the “stable-diffusion-1-5” model, “s-d-2-1” represents the outcomes of the “stable-diffusion-2-1” model, and “Our” indicates results generated by our PrototypeAgent.
7.2. User Study 2: Usage Study
In order to obtain insightful feedback on our tool, we compared our PrototypeAgent with emerging design tools; we conducted another user study followed by expert interviews. This study involved a comparative analysis using our tool, Vercel’s V0(Vercel, 2024)—a UI code-based design tool, and Uizard(Uizard, 2024)—a Prompt to UI design tool. After engaging with these tools, participants completed a 5-point Likert Scale questionnaire based on 5 questions. The questionnaire aimed to assess how well our system addressed five design goals identified in Section 3, in terms of all aspects of performance and interactiveness.
7.2.1. Procedure
The study began with an introductory session to acquaint participants with the study procedure. Participants were then asked to create 2 GUI prototypes using our tool, Vercel’s V0, and Uizard, based on provided design purposes and corresponding wireframes. Those 2 prototype design purposes are random obtained from 20 different design purposes, which are selected from the Screen2Word test dataset for GUI generation, with each containing an average of 5 components. Participants were allowed to modify the input description until they were satisfied with the generated results, and the number of manual modifications made by users was recorded.
Following task completion, we conducted a semi-structured survey with each participant. This survey included 5 main questions on a 5-point Likert Scale, focusing on eliciting feedback about our tool’s performance relative to the five identified design findings. Additionally, participants were asked for suggestions on potential improvements. The survey questions were as follows:
  1. (1)
    How effective was the system in providing relevant design suggestions or resources from the knowledge base? (Related to F1: Knowledge Base Effectiveness)
  2. (2)
    Was the input method intuitive, and did it allow you to express the design intent clearly and precisely? (Related to F2: Interaction Modality)
  3. (3)
    Did the tool provide useful suggestions to clarify your design intent? (Related to F3: Implicit Intent Clarification)
  4. (4)
    Were the intermediate results useful in aligning the design with your intent, and did they enhance interaction? (Related to F4: Human-Centered Intent-Design Alignment)
  5. (5)
    Did the tool maintain consistent semantics and style across all design elements, ensuring a cohesive output? (Related to F5: Ensuring Design Coherence)
Figure 9. Results of the usage study. Blue represents Vercel’s V0, green denotes Uizard, and red indicates our system PrototypeAgent.
7.2.2. Results & Discussion
The feedback gathered from our semi-structured survey offered critical insights into the performance of our tool in relation to the challenges outlined in Section 3, and highlighted potential areas for enhancement. Fig. 9 shows the results.
Knowledge Base Effectiveness (F1). Participants rated the Relevant Designs, with Uizard achieving a score of 3.5, our tool closely following at 3.4, and Vercel’s V0 receiving 3.0. In follow-up interviews, we found that 9/16 participants believed our tool excelled in generating detailed elements, such as text, icons, and images. This feedback highlights the strength of our decoupled generation method, which maximizes the system’s ability to create detailed, functional design elements. On the other hand, 7/16 participants appreciated Uizard’s overall style design, describing it as visually more appealing. In addition, 4/16 participants praised our system’s flexibility, particularly the ability to import and integrate their own data into the generation process, allowing for customization. They suggested that the effectiveness of this feature could be verified in the future.
This feedback highlights the potential of our tool in integrating design libraries—a feature lacking in Vercel’s V0 and partially implemented in Uizard. Our PrototypeAgent offers a more customizable and functional design experience.
Interaction Modality (F2). Our system led in reflecting design intent with a score of 4.1, compared to Uizard’s 3.6 and Vercel’s 2.5. Participants valued our tool’s use of both wireframes and prompts, which Uizard lacks, and found Vercel’s code generation to not adhere closely to layout sizes. A participant remarked, “PrototypeAgent’s method of rendering components from wireframes and prompts is impressive. The design of layout parts encompasses functional design and is labor-intensive. AI assistance in this area is much needed.”
Implicit Intent Clarification (F3). Our PrototypeAgent scored 4.0 for Implicit Intent Clarification, outperforming Uizard (3.2) and Vercel’s V0 (1.6). 14/16 participants found our clarification ability particularly helpful, especially for themes and app categories. One participant noted, ”The automatic clarification is more professional than my own input, saving me from guessing what the AI needs.” This feature allowed designers to focus more on achieving their desired results, reducing guesswork and effort. On average, designers using our tool made 2.6 revisions to reach their desired design. In comparison, Uizard required 4.3 revisions, and Vercel’s V0 required 6.1 revisions, with many users still unsatisfied, citing lack of control over the generated code format.
Regarding color clarification, designers with more than three years of experience expressed the need for tools that respect company design constraints, particularly for theme colors. They suggested that future improvements could include interactive visualizations with customizable color palettes to enable more precise adjustments.
Human-Centered Intent-Design Alignment(F4). Our tool was highly regarded for its intermediate generation process in helping with human-in-the-loop intent-design alignment, achieving a score of 3.8. In comparison, Uizard scored 3.5, while Vercel’s V0 lagged behind with a 2.5. Notably, 14 out of 16 participants found the intermediate steps in our tool exceedingly beneficial for interactive engagement with the AI. A participant shared a compelling example: “While working on a job search page, I was impressed by the UI designs produced by the tool. They were professional and included all necessary details, such as job categories, locations, and salaries, making it feel like a real, complete design.” This feedback underscores our tool’s proficiency in providing clear, informative intermediate steps that facilitate effective user interaction and ensure the generation of high-quality prototypes.
Participants highlighted a distinctive advantage of our tool over others: while Vercel’s V0 allows code-based component edits and Uizard enables manual adjustments, our tool uniquely presents both the components and critical information during their generation, including prompts. This feature was particularly lauded for its precision and convenience. One participant expressed, “I was surprised and pleased by the direct control over fine-tuning local components through prompts. This is a stark contrast to Vercel’s V0, which often struggles with accurately locating specific components for adjustments.” Echoing this sentiment, 15 out of 16 participants appreciated our tool’s top-down editing support, which they found met their needs from broader adjustments to more detailed refinements, significantly aiding them in realizing their design vision accurately.
Ensuring Design Coherence (F5). In the critical aspect of maintaining design coherence, our tool and Uizard both achieved a high score of 4.2, clearly outperforming Vercel’s V0’s 3.5. A majority of the participants, 10 out of 16, remarked on the effective maintenance of colour and image style consistency by both our tool and Uizard. They noted that these tools produced designs with cohesive colour schemes and consistent visual elements.
However, while Vercel V0’s outputs were generally competent, its lack of image components and tendency towards simplistic black and white colour schemes were seen as areas needing improvement. Beyond the aesthetic aspects of design consistency, participants also highlighted the importance of colour accessibility. A participant pointed out a significant oversight in current tools, stating, “Although the tools offer impressive theme colour designs, a key shortfall is their lack of consideration for colour accessibility, making some designs impractical for real-world application.” This feedback underscores the necessity for future tools to incorporate colour accessibility as a fundamental component of design consistency.
Weaknesses and improvements. Participants offered valuable feedback on potential improvements and identified exciting directions for future enhancements. Several participants saw great potential in augmenting the user experience by introducing automated wireframe generation from high-level descriptions. This feature could significantly streamline the design process by reducing manual steps, thereby increasing efficiency. A participant highlighted this by saying, “It would be really helpful if the tool could auto-generate wireframes based on high-level descriptions. It would save me more time.”
Addressing specific failure cases, one participant pointed out a drawback in the current image generation process: “Sometimes the generated images for small, simple areas are overly complex. Simplified vector images like icons might be more effective.” This observation led to the suggestion that the tool should differentiate between standard images and vector graphics in future iterations for improved functionality.
Furthermore, in terms of usability and accessibility, participants suggested that aligning the tool more closely with established UI design guidelines, such as Material Design, would greatly enhance its utility. One participant emphasized the importance of functionality alongside aesthetics: “Enhancing the tool to adhere to UI design guidelines, such as Material Design, would significantly improve its usability. A prototype should not only look good but should also be interactive and user-friendly.” Building upon this, another added, “Creating visually stunning interfaces is great, but incorporating accessibility guidelines is crucial for making designs truly universal.”
Answer to RQ4: The usage study underscored the effectiveness of our tool in refining the UI prototyping process. It adeptly balances automation with customization and offers user-friendly interactions. Particularly notable is its capability to better communicate design intent and provide transparent, explainable AI-assisted intermediate results. This presentation of both components and vital information during generation has been highly praised for its precision and convenience, facilitating designer workflows more effectively than other emerging design tools.
8. Discussion and Future Work
Our current work, PrototypeAgent, focuses on improving the design generation process and enhancing the quality of prototypes. It draws inspiration from a rich knowledge base and primarily relies on high-level design descriptions and wireframe layouts to translate design intent into tangible design prototypes. Looking ahead, there are several future directions for improvements and extensions:
1) Exploring Enhanced AI-Human Collaboration in UI Design A key future direction involves refining input modalities to support a broader range of intent-guided inputs, such as sketching wireframes or referencing design guidelines. The aim is not merely to expand the input options but to minimize user effort while helping designers express their intent more precisely with less input. For instance, our current system’s Implicit Intent Clarification feature already reduces user workload by automatically clarifying and refining design intent.
Additionally, we can explore how AI can further assist designers by automating more aspects of the process, such as generating wireframes automatically, thereby reducing manual effort. Although it would be relatively straightforward to remove the wireframe creation input and allow our system to automatically retrieve and generate wireframes, our current focus is on achieving accurate input expression. For this reason, we chose not to include this feature at this stage. However, we recognize the potential benefits of automated wireframe generation and see it as a valuable direction for future enhancements.
2) Usability and Accessibility Enhancements. As we continue to improve PrototypeAgent, we are committed to augmenting our system with standard mobile device user interface guidelines like material design (Google, 2023). These improvements will further enhance our system’s usability and accessibility, ensuring our prototypes not only look appealing but also offer intuitive and user-friendly interactions. To further improve the design process, we could incorporate usability testing and visual reports that highlight potential issues early in the design stage, even during the generation process. Early detection and correction of design flaws will help reduce the cost of revisions, making the system more efficient and accessible.
3) Contextual Relationships and Multi-page Designs. Recognizing that UI designs often span multiple pages, in the future, we could discuss and expand on contextual relationships in our design generation. This will allow us to produce designs for multiple pages simultaneously in future iterations, ensuring a cohesive style across these pages and thereby enhancing efficiency.
4) Tool Integration and Design-to-Code. With an aim to seamlessly incorporate our system into designers’ workflow, we imagine PrototypeAgent could be developed as a plugin that integrates smoothly with popular wireframe sketching software, such as Figma (Figma, 2024) and Sketch (Sketch, 2024). This would widen our system’s functionality by providing designers with robust editing capabilities directly within these platforms. Additionally, future work could also explore ways to automate design-to-code conversion to provide designers and developers with an end-to-end solution.
9. Conclusion
In this paper, we identified five key findings in designers’ normal workflow with AI design tools. To address these existing gaps, we proposed PrototypeAgent, a novel multi-agent prototype generation system, delicately balancing automation with customization. Given human-provided descriptions and wireframe layouts, our system iteratively refines these inputs into engaging, high-fidelity design prototypes, maintaining aesthetic harmony and aligning with the design intent. Beyond generation, PrototypeAgent not only automates Implicit Intent Clarification for designers but also provides explainable intermediate results to enhance Human-Centered Intent-Design Alignment. Our quantitative and qualitative evaluations further corroborate the potential of our approach to significantly improve the UI / UX design process. Going forward, we will continue improving our work, like enabling automated wireframe generation and supporting dynamic component integration for a more universal, user-friendly, efficient and creative design process.